log3x^2=2log3 4 - 4 log 3 5
Answer by ewatrrr(24785)

 
Hi,
Note:






log3x^2=2log3 4 - 4 log 3 5


x^2(5^4)/16 = 1
x^2 = 16/5^4
x = ± 4/25